# ANPrompt: Anti-noise Prompt Tuning for Vision-Language Models

### v2

@@ Abstract
Prompt tuning has emerged as an efficient and effective technique for adapting vision-language models (VLMs) with low computational overhead. However, existing methods often overlook the vulnerability of prompt-tuned VLMs to weak semantic perturbations—such as subtle image or text noise—that degrade their generalization to unseen classes. To address this limitation, we propose ANPrompt, a novel prompt tuning framework designed to enhance robustness under such perturbations. ANPrompt first constructs weak noise text features by fusing original and noise-perturbed text embeddings, which are then clustered to form noise prompts. These noise prompts are integrated with learnable prompt tokens to generate anti-noise prompts, which are injected into the deeper layers of both image and text encoders. To further capture the noise-aware visual semantics, ANPrompt computes the Noise-Resistant Visual Prompt Prototype (NRVPP) by averaging the output prompt tokens from the vision encoder. Finally, ANPrompt introduces alignment, robustness, and anti-noise objectives by computing a Weak semantic noise Alignment Loss (WALoss) alongside the standard cross-entropy and sim loss. Experiments across 11 benchmarks demonstrate that ANPrompt consistently outperforms existing prompt tuning approaches, achieving superior robustness to semantic noise and improved generalization to novel categories.
Introduction
Vision-language models (VLMs) like CLIP have demonstrated impressive performance on open-vocabulary recognition tasks  (Hua et al. 2025; Weng et al. 2023; Xu et al. 2024), but adapting such large models to downstream tasks via full fine-tuning can be computationally prohibitive. To address this, prompt tuning (Zhu et al. 2024; Lafon et al. 2024; Shi et al. 2024) has emerged as a lightweight alternative by optimizing only a small set of learnable tokens, achieving state-of-the-art results across various domains. Despite its efficiency, early prompt tuning methods such as CoOp (Zhou et al. 2022b) often overfit to seen classes, resulting in limited generalization to novel categories and vulnerability to subtle semantic perturbations.
To enhance transferability, recent approaches have integrated inductive biases into the prompting process. For instance, CoCoOp conditions prompts on image features to mitigate overfitting; KgCoOp constrains semantic drift by leveraging frozen text embeddings; and MaPLe adopts multimodal prompts to achieve more robust alignment. More advanced techniques such as PromptSRC and CoPrompt regularize prompt learning using frozen zero-shot logits, while MMRL and TAC construct shared or clustered semantic spaces. Although designed to reduce overfitting, consistency constraints based on frozen features may introduce noise information such as misleading text or visual clutter caused by domain gaps, ultimately weakening the robustness of Vision-Language Models.
@@ Introduction
Figure 1: Weak semantic perturbations in image and text: CAMs may highlight irrelevant regions (e.g., background or co-occurring objects), while subtle textual cues (e.g., “dark stripes”) may mislead class semantics (e.g., shift from “cat” to “tiger”).
Figure 2: Different with existing methods use frozen logits as constraint, risking overfitting under weak semantic perturbations. ANPrompt introduces weak semantic perturbations to generate soft logits and anti-prompt to improves generalization.
As shown in Figure 1, we define subtle noise in images and texts as weak semantic perturbations, which appear as either misleading text (e.g., “a cat resembling a small tiger”) or visual noise (e.g., background clutter or co-occurring objects). While high-quality image-text pairs exhibit clear object focus and accurate descriptions, existing prompt tuning methods often struggle to handle such subtle shifts, leading to significant performance drops when adapting to novel categories. This highlights the urgent need for VLMs capable of learning robust, noise-resilient cross-modal representations to maintain alignment consistency and improve generalization.
To address the issue of weak semantic perturbations, we propose ANPrompt — a prompt tuning framework specifically designed to enhance robustness under such perturbations. As illustrated in Figure 2, existing methods rely on frozen feature constraints and thus tend to suffer from overfitting when exposed to weak semantic perturbations. In contrast, our ANPrompt framework intentionally incorporates weak semantic perturbation information into prompt tokens and logits, thereby boosting the robustness of both the learned prompts and class distributions in scenarios involving weak semantic perturbations. Specifically, ANPrompt randomly samples two intra-class descriptions from a language cache: a main text that conveys core semantics, and a noise text that introduces minor linguistic variations. Their fusion generates a Weak Noise Frozen Text Feature, which incorporates weak semantic perturbation information while preserving class identity. These features are clustered into noise prompts and injected as biases into learnable tokens, forming Anti-noise Prompts. These prompts are then inserted into the higher layers of both text and vision encoders to introduce weak semantic perturbation information at the prompt representation level. After encoding, ANPrompt extracts image Feature, text Feature, and the Noise-Resistant Visual Prompt Prototype (NRVPP). The latter is obtained by mean-pooling the visual prompt tokens from the final layer of the visual encoder. Since Anti-noise Prompts are injected into deeper encoder layers, the resulting NRVPP inherit weak semantic perturbations.
To stabilize training under both weak semantic perturbations, we introduce a KL divergence-based Weak Alignment Loss (WALoss) between Weak Noise Logits and Logits set, which enforces alignment consistency under controlled weak semantic perturbations. Additionally, we employ a CE and sim loss to enable the model to learn task-relevant knowledge under weak semantic perturbations. By combining these losses, ANPrompt can improve the robustness of VLMs under weak semantic perturbations.
Extensive experiments conducted on 11 benchmarks and base-to-new splits show that ANPrompt significantly outperforms existing prompt tuning methods, achieving state-of-the-art performance in terms of robustness and generalization.
@@
  • We propose ANPrompt, a prompt tuning framework that improve the robustness of VLMs by injecting weak semantic perturbations into both prompt tokens and logits pathways.
  • We introduce a Weak semantic noise Alignment Loss (WALoss) to enforce alignment consistency under controlled weak semantic perturbations.
  • We develop a weak semantic perturbations prompt injection mechanism that clusters weak noise frozen text features into noise prompts and injects them into learnable tokens to get anti-noise prompt, which is used to inject weak semantic perturbations at the prompt representation level.
  • ANPrompt achieves state-of-the-art performance on 11 benchmark datasets under base-to-new generalization settings.
Related Work
Vision-Language Models
@@
Pretrained vision-language models (VLMs) such as CLIP (Radford et al. 2021) and ALIGN (Jia et al. 2021) achieve strong open-vocabulary and cross-modal capabilities through contrastive learning on large-scale image-text pairs. Building on this foundation, models like BLIP (Li et al. 2022), FILIP (Yao et al. 2021), Florence (Yuan et al. 2021), and LiT (Zhai et al. 2022) enhance fine-grained alignment and data efficiency via retrieval augmentation(Li et al. 2024a), self-supervised learning, and stronger backbones. NLIP (Huang et al. 2023) further improves robustness by introducing noise-harmonization based on alignment uncertainty and noise-completion via concept-guided caption generation. These advances enable effective zero-shot and few-shot transfer for tasks such as retrieval, classification, and grounding. However, adapting such large models to downstream tasks (Li et al. 2025a, 2023) remains costly due to domain shifts and the high overhead of full fine-tuning. To address this, parameter-efficient tuning methods—such as adapter-based approaches (Gao et al. 2024; Yang et al. 2024b; Song et al. 2023) and prompt learning (Zhou et al. 2022b, a; Khattak et al. 2024)—have been proposed to enable lightweight adaptation without retraining the entire model.
Prompt Learning
@@
Prompt learning has become an efficient strategy to adapt VLMs without full fine-tuning. Early methods such as CoOp (Zhou et al. 2022b) and CoCoOp (Zhou et al. 2022a) explore soft and image-conditioned prompts to enhance generalization. Subsequent works improve robustness and transferability through multimodal regularization (MaPLe (Khattak et al. 2023a), PromptSRC (Khattak et al. 2023b)), gradient alignment (ProGrad (Zhu et al. 2023)), hierarchical and knowledge-guided prompts (HPT (Wang et al. 2024), TCP (Yao, Zhang, and Xu 2024)), and distillation-based initialization (CLIP-kd (Yang et al. 2024a), ComKD-CLIP (Chen et al. 2024), PromptKD (Li et al. 2024b), CasPL (Wu et al. 2024)). Other methods focus on attribution-based prompting (ATPrompt (Li et al. 2024c)), two-stage decoupled optimization (DPC (Li et al. 2025b), DePT (Zhang et al. 2024), 2SFS (Farina et al. 2025)), or cross-modal adapters (MMA (Yang et al. 2024b)) for few-shot settings. Consistency learning (CoPrompt (Roy and Etemad 2024)) and task-aware clustering (TAC (Hao et al. 2025)) further enhance adaptability. However, these approaches largely target generalization and efficiency, with limited attention to robustness against weak semantic perturbations.
Prompting under Weak Semantic Perturbation Noise
@@
Recent works have explored robust prompt learning under noisy conditions. Methods such as NLPrompt (Pan et al. 2025), JoAPR (Guo and Gu 2024), and ArGue (Tian et al. 2024) mitigate label noise through loss design, sample filtering, and LLM-generated attributes, while ClipCleaner (Feng, Tzimiropoulos, and Patras 2024) leverages CLIP-based sample selection to decouple training from noise filtering. However, these approaches mainly focus on filtering noise, which risks discarding informative variations and often limits generalization. In contrast, we propose ANPrompt, which embraces weak semantic perturbations as supervision signals. By integrating perturbed prompts with clustering-based bias representations, ANPrompt enables robust and discriminative prompt learning without relying on brittle filtering heuristics.
Figure 3: ANPrompt randomly samples a main and a noise sentence from the same class to construct the Weak Noise Frozen Text Feature, which is used to generate Anti-noise Prompts. These prompts are injected into the encoder’s deep layers. During encoding, we extract the Prompted Image/Text Features and the Noise-Resistant Visual Prompt Prototype (NRVPP) to compute four types of logits, which are supervised by WALoss and auxiliary losses to improve robustness against weak semantic perturbations.
Methodology
Preliminary
@@
Vision-language models (VLMs) such as CLIP (Radford et al. 2021) learn a joint embedding space for images and texts through contrastive learning over large-scale image-text pairs. These models exhibit strong zero-shot generalization but are highly sensitive to semantic variations in either modality.
@@
A typical VLM consists of a vision encoder and a text encoder. The vision encoder tokenizes the input image into patches, encodes them through Transformer layers, and outputs a class token representing global semantics. The text encoder processes class-dependent prompts (e.g., “a photo of a [CLASS]”), encodes them with Transformers, and uses the [EOS] token as the prompt embedding.
@@
Overview
We propose ANPrompt — a robust prompt tuning framework that injects weak semantic perturbations into the prompting and alignment processes to enhance the robustness of VLMs under such perturbations. Specifically, ANPrompt simulates weak semantic perturbations by fusing two intra-class textual descriptions from the language cache, generating a Weak Noise Frozen Text Feature. These Weak Noise Frozen Text Features are clustered to obtain noise prompts, which are then projected and added to learnable prompt tokens to form Anti-noise Prompts. These prompts are inserted into the deeper layers of both vision and text encoders, enabling ANPrompt to inject weak semantic perturbations at the prompt level. Subsequently, prompted image features, prompted text features, and NRVPP can be obtained for computing alignment logits, robustness logits, weak noise logits, and anti-noise logits. During training, ANPrompt combines the Weak Semantic Noise Alignment Loss (WALoss), cross-entropy loss, and SimLoss to enhance the robustness of VLMs under weak semantic perturbations.
ANPrompt Framework
@@
As shown in Figure 3. The ANPrompt framework proceeds through four sequential stages: (1) Weak Noise Frozen Text Feature Construction, (2)Anti-noise Prompt Generation, (3) Feature Encoding, and (4) Logit Computation.

### v3

# Robust Prompt Tuning for Vision-Language Models with Mild Semantic Noise

@@ Abstract
Prompt tuning has shown promising results, but its robustness and generalization to unseen categories remain limited. Through our experiments, we demonstrate that the complete removal of semantic noise is a key factor restricting robustness. Existing methods typically suppress or filter out semantic noise in the prompt space, inadvertently hindering the model’s robustness and its ability to generalize to unseen categories. To address this, we propose ANPrompt, a robust prompt tuning framework that actively incorporates weak semantic noise. By clustering weakly perturbed features into noise prompts and integrating them with learnable tokens in both the text and vision encoders, ANPrompt ensures controlled exposure to semantic variations. To enhance the visual pathway, we introduce the Noise-Resistant Visual Prompt Prototype (NRVPP), which stabilizes visual semantics under weak perturbations. Additionally, we propose a Weak Alignment Loss (WALoss) at the logits level to enforce consistency between clean and perturbed predictions, providing stable supervision. By combining weak semantic noise exposure with logits-based consistency, ANPrompt prevents overfitting to specific phrasings while preserving semantic integrity. Extensive experiments across 11 benchmarks, including base-to-new splits, show that ANPrompt consistently outperforms existing prompt tuning methods, offering superior robustness to semantic noise and improved generalization across tasks.
Figure 1: This figure compares the accuracy of various methods, including DAPT, Synreplace, Mask, Shuffle, and Drop. The results show that incorporating weak semantic noise achieves superior performance, outperforming DAPT (79.35%) and other baseline methods. This highlights the effectiveness of weak semantic noise injection in enhancing model robustness and generalization.
Figure 2: Framework comparison with representative filtering-based methods such as DAPT and ArGue. While these approaches constrain frozen logits and risk overfitting under weak semantic perturbations, the proposed ANPrompt leverages weak semantic perturbations to generate soft logits and anti-noise prompts, thereby improving generalization.
@@
Figure 3: Illustration of weak semantic perturbations in both image and text. Class activation maps (CAMs) may highlight irrelevant regions such as background or co-occurring objects, while subtle textual cues (e.g., “dark stripes”) can shift class semantics from “cat” to “tiger”. These examples demonstrate how weak semantic noise can mislead recognition and motivate the need for robust prompt learning.
Figure 4: Quantitative comparison of different noise strategies in terms of Text Shift (TS), Logit Preservation Rate (LPR), and Accuracy Shift (AS). The results show that strong perturbations (drop, mask, shuffle, synonym replacement) cause semantic distortion and performance degradation, whereas our weak semantic perturbations (WSP) maintain negligible shift, high prediction consistency, and nearly zero accuracy loss.
I Introduction
Vision-language models (VLMs), such as CLIP, have demonstrated remarkable performance on open-vocabulary recognition tasks openworldauc ; openvclip ; xu2024llava . However, adapting these large-scale models to downstream tasks is often infeasible due to the high computational cost of full fine-tuning. Prompt tuning awt ; gallop ; qnet has emerged as an efficient alternative, updating only a small set of learnable tokens. Early approaches, such as CoOp coop , tend to overfit to seen classes, leading to poor generalization to novel categories and sensitivity to semantic perturbations. Subsequent methods, including CoCoOp cocoop , KgCoOp kgcoop , MaPLe maple , and MMRL mmrl , introduce inductive biases or shared semantic spaces to enhance robustness, but they still struggle under noisy conditions.
Existing methods for handling noisy information in prompt tuning typically follow a filtering philosophy, where noise is suppressed or eliminated. Specifically, methods like DAPT dapt and ArGueargue align foreground features with textual descriptions while suppressing irrelevant background, as shown in Figure 2(a). However, these techniques often risk discarding valuable semantic variations and struggle to adapt to subtle, real-world perturbations. Furthermore, we demonstrate that filtering or suppressing noise, as done by methods like DAPT (79.35%) and Synreplace (81.44%), often limits robustness. These approaches suppress irrelevant background noise, which not only discards valuable semantic information but also hinders the model’s generalization and robustness to diverse real-world perturbations.
In contrast, our method actively integrates weak semantic noise, resulting in a notable increase in performance. As shown in Figure 1, methods that filter or suppress noise show limited improvement, while our approach of incorporating weak semantic noise consistently outperforms these baseline methods. This emphasizes that actively incorporating weak semantic noise is more effective than filtering it, enabling the model to generalize better across unseen categories and enhancing its robustness to perturbations. As shown in Figure 2(b), our method actively injects weak semantic noise, enhancing robustness by integrating it with constraint-based techniques, rather than suppressing it. This allows the model to better adapt to diverse inputs without sacrificing semantic integrity.
Noise injection has also been explored in the field of multimodal representation learning. Related works, such as noise1 , have proposed several text-based noise injection strategies, including strong perturbations like token dropping, masking, word shuffling, and synonym replacement, as illustrated in Figure 3. While these strong perturbations introduce variation, they often distort the original semantics and degrade model generalization, as demonstrated in the top-left part of the figure. However, we find that such strong noise injection methods are not well-suited for the Prompt tuning domain. To address this issue, we propose a more gradual noise injection approach, introducing weak semantic perturbations, which apply subtle shifts in both the image and text domains. By incorporating a logits constraint, shown in the bottom portion of Figure 3, our method ensures these perturbations preserve class identity while enhancing robustness. The stable Noise ZeroShot Logits generated by this approach help the model generalize better to unseen variations without the disruptive effects of strong noise perturbations.
To substantiate this, we run a controlled comparison under the base-to-new setting, as shown in Figure 4. Strong perturbations severely distort semantics: token dropping yields $\mathrm{TS}=0.0196$ and $\mathrm{AS}=0.618$, masking further increases distortion with $\mathrm{TS}=0.0300$ and $\mathrm{AS}=0.923$, and word shuffling reaches $\mathrm{TS}=0.0573$ with $\mathrm{AS}=0.438$. Even synonym replacement introduces instability ($\mathrm{TS}=0.0090$, $\mathrm{AS}=0.00013$). In contrast, weak semantic noise preserves class identity, exhibiting negligible text shift ($\mathrm{TS}=6.3\times 10^{-6}$ at $\epsilon=0.01$), high prediction consistency ($\mathrm{LPR}\approx 0.98$), and nearly zero accuracy degradation. These results suggest that controlled weak noise exposure can enhance robustness without sacrificing semantic fidelity.
Building on these insights, we propose ANPrompt, a prompt-tuning framework that actively leverages weak semantic noise rather than suppressing it. Specifically, ANPrompt clusters weak noise features into representative prompts and integrates them with learnable tokens in both text and vision encoders. As shown in Figure 2(b), our approach contrasts with previous methods that focus on noise suppression. By actively incorporating weak semantic noise, the model improves its robustness to unseen categories, while effectively enhancing generalization and maintaining semantic consistency. To further capture noise-aware visual semantics, we introduce a Noise-Resistant Visual Prompt Prototype (NRVPP), which provides stable visual representations under weak perturbations. This controlled noise exposure regularizes prompt representations, mitigates overfitting to specific textual expressions, and promotes learning of more robust class-level semantics. Furthermore, we introduce a Weak Alignment Loss (WALoss) that enforces consistency between clean and perturbed predictions. Unlike feature-level perturbations, which may disrupt intermediate representations and lead to instability, enforcing consistency at the logits level provides a more stable supervision signal. This directly regularizes the decision boundary and improves robustness without compromising the model’s discriminative capacity. The benefits of WALoss align with our overall approach of controlled noise exposure, further enhancing the model’s ability to adapt to real-world variations without losing accuracy or generalization.
@@ I Introduction
  • We propose ANPrompt, a prompt tuning framework that improves robustness by incorporating weak semantic noise into both prompt tokens and logits.
  • We introduce a Weak Alignment Loss (WALoss) that enforces prediction consistency at the logits level, providing stable supervision and stronger generalization compared to feature-level perturbations.
  • We exploit weak semantic noise as a source of natural linguistic diversity, preserving class identity while preventing overfitting to specific phrasings, thereby enhancing base-to-new generalization compared with strong noise strategies.
  • We conduct extensive experiments on 11 datasets and base-to-new splits, demonstrating that ANPrompt achieves state-of-the-art robustness and generalization compared with existing prompt tuning methods.
II Related Works
II-A Vision-Language Models
@@
Pretrained vision-language models (VLMs)l1 ; l2 ; l3 ; l4 ; l5 such as CLIP clip and ALIGN align achieve strong open-vocabulary and cross-modal capabilities through contrastive learning on large-scale image-text pairs. Building on this foundation, models like BLIP blip , FILIP filip , Florence florence , and LiT lit enhance fine-grained alignment and data efficiency via retrieval augmentationli2024freestyleret , self-supervised learning, and stronger backbones. NLIP NLIP further improves robustness by introducing noise-harmonization based on alignment uncertainty and noise-completion via concept-guided caption generation. These advances enable effective zero-shot and few-shot transfer for tasks such as retrieval, classification, and grounding. However, adapting such large models to downstream tasks li2025beyond ; li2023weakly ; pcl ; hipl ; sod remains costly due to domain shifts and the high overhead of full fine-tuning. To address this, parameter-efficient tuning methods—such as adapter-based approaches clipadapter ; mma ; metaadapter and prompt learning protext ; cpr ; pmp —have been proposed to enable lightweight adaptation without retraining the entire model.
II-B Prompt Learning
@@
Prompt learning has become an efficient strategy to adapt VLMs without full fine-tuning. Early methods such as CoOp coop and CoCoOp cocoop explore soft and image-conditioned prompts to enhance generalization. Subsequent works improve robustness and transferability through multimodal regularization (MaPLe maple , PromptSRC promptsrc ), gradient alignment (ProGrad prograd ), hierarchical and knowledge-guided prompts (HPT hpt , TCP tcp ), and distillation-based initialization (CLIP-kd clipkd , ComKD-CLIP comkdclip , PromptKD promptkd , CasPL caspl ). Other methods focus on attribution-based prompting (ATPrompt atprompt ), two-stage decoupled optimization (DPC dpc , DePT dept , 2SFS 2sfs ), or cross-modal adapters (MMA mma ) for few-shot settings. Consistency learning (CoPrompt coprompt ) and task-aware clustering (TAC tac ) further enhance adaptability. However, these approaches largely target generalization and efficiency, with limited attention to robustness against weak semantic perturbations.
II-C Prompting under Weak Semantic Perturbation Noise
@@
Recent works have explored robust prompt learning under noisy conditions. Methods such as NLPrompt nlprompt , JoAPR joapr , and ArGue argue mitigate label noise through loss design, sample filtering, and LLM-generated attributes, while ClipCleaner clipcleaner leverages CLIP-based sample selection to decouple training from noise filtering. However, these approaches mainly focus on filtering noise, which risks discarding informative variations and often limits generalization. In contrast, we propose ANPrompt, which embraces weak semantic perturbations as supervision signals. By integrating perturbed prompts with clustering-based bias representations, ANPrompt enables robust and discriminative prompt learning without relying on brittle filtering heuristics.
III Method
III-A Preliminaries
@@
Vision-language models (VLMs) such as CLIP clip learn a joint embedding space for images and texts through contrastive learning over large-scale image-text pairs. These models exhibit strong zero-shot generalization but are highly sensitive to semantic variations in either modality.
@@
A typical VLM consists of a vision encoder and a text encoder. The vision encoder tokenizes an input image into patches, encodes them through Transformer layers, and outputs a class token representing global semantics. The text encoder processes class-dependent prompts (e.g., “a photo of a [CLASS]”), encodes them with Transformers, and uses the [EOS] token as the prompt embedding.
@@
Figure 5: Overview of ANPrompt. For each class, a main and a noise sentence are sampled to construct the Weak Noise Frozen Text Feature, which is clustered into noise prompts. These are combined with learnable tokens to generate Anti-noise Prompts, which are injected into the last two layers of both vision and text encoders. During encoding, we obtain Prompted Image/Text Features and the Noise-Resistant Visual Prompt Prototype (NRVPP) to compute four types of logits, supervised by WALoss and auxiliary losses to improve robustness against weak semantic perturbations.
III-B ANPrompt Framework
We propose ANPrompt, a robust prompt tuning framework that explicitly leverages weak semantic noise to enhance the robustness of VLMs. Specifically, ANPrompt simulates weak perturbations by fusing intra-class textual descriptions, generating Weak Noise Frozen Text Features. These features are clustered to form noise prompts, which are added to learnable prompt tokens to construct Anti-noise Prompts. The Anti-noise Prompts are injected into the last two transformer blocks of both vision and text encoders by prepending them as additional tokens, thereby exposing the model to weak semantic perturbations during prompt learning. After encoding, we obtain prompted image features, prompted text features, and the Noise-Resistant Visual Prompt Prototype (NRVPP), which are used to compute four types of logits: alignment, robustness, weak noise, and anti-noise logits. During training, ANPrompt combines the Weak Alignment Loss (WALoss) with cross-entropy and similarity losses to enhance robustness against weak semantic perturbations. Specifically, WALoss encourages the model to maintain prediction consistency between clean and perturbed inputs at the logits level, ensuring that the model remains stable even when exposed to subtle changes in the data. By incorporating weak semantic noise, ANPrompt regularizes the decision boundary, making it more resilient to small semantic shifts without sacrificing accuracy. Unlike previous methods, which primarily employ generic prompt regularization based on cosine similarity or feature-level perturbations, ANPrompt takes a more explicit approach by integrating weak semantic noise directly into prompt tokens. It enforces consistency at the logits level through WALoss, which provides a more stable supervision signal compared to traditional methods. This mechanism is distinctive in its ability to simultaneously preserve class identity and improve generalization, making ANPrompt a more effective solution for handling noise in vision-language models.
@@ III-B ANPrompt Framework
As illustrated in Figure 5, the ANPrompt framework proceeds through four sequential stages: (1) Weak Noise Frozen Text Feature Generation, (2) Anti-noise Prompt Construction, (3) Feature Encoding, and (4) Logit Computation.
III-B1 Weak Noise Frozen Text Feature Generation